# roundcorner=10pt, linewidth=1pt, innertopmargin=10pt, innerbottommargin=10pt, innerrightmargin=10pt, innerleftmargin=10pt, backgroundcolor=gray!10 Agent Hospital: A Simulacrum of Hospital with Evolvable Medical Agents

### v1

@@ 5.2. Experimental Results
From the experimental results, we have the following conclusions. Firstly, the proposed MedAgent-Zero strategy effectively enhances doctor Agents on the three tasks, with the cumulative accuracy on 10,000 training samples showing a continuous increase. The best performances of examination, diagnosis, and treatment are 88%, 95.6%, and 77.6%, respectively. It shows our agent evolves during the training phase, just as human doctors become experienced after treating thousands of patients. Furthermore, agent evolution is more efficient than human, as human doctors may take over two years to treat ten thousand patients.
@@ 5.3.3. Case Study
To further analyze the usefulness of the medical record library and experience base, we conduct a case study here. Table 1 illustrates the performance of the experience base, medical record library separately, and MedAgent-Zero on the three tasks of a patient case. At first, all answers are incorrect without experience or medical records. When only the experience is added, the examination and diagnosis responses are correct, but the treatment response is incorrect. When only the relevant medical records are added, the examination response is incorrect, while diagnosis and treatment level responses are correct. Finally, when both the experience and records are added simultaneously, the doctor can provide correct responses for all three tasks.
@@
| Strategy | Examination | Diagnosis | Treatment |
| Only Experience Base | |||
| Only Medical Records | |||
| MedAgent-Zero |
@@ Appendix A Contributions of Authors
Appendix B Involved Disease Knowledge
B.1. Diseases
@@ B.2. Data Distribution
B.3. An Example of electronic Health Records of a Patient

### v2

@@ 5.2. Experimental Results
From the experimental results, we have the following conclusions. Firstly, the proposed MedAgent-Zero strategy effectively enhances doctor Agents on the three tasks, with the cumulative accuracy on 10,000 training samples showing a continuous increase. The best performances of examination, diagnosis, and treatment are 88%, 96.6%, and 81%, respectively. It shows our agent evolves during the training phase, just as human doctors become experienced after treating thousands of patients. Furthermore, agent evolution is more efficient than human, as human doctors may take over two years to treat ten thousand patients.
@@ 5.3.3. Case Study
To further analyze the usefulness of the medical record library and experience base, we conduct a case study here. Table LABEL:case_table illustrates the performance of the experience base, medical record library separately, and MedAgent-Zero on the three tasks of a patient case. At first, all answers are incorrect without experience or medical records. When only the experience is added, the examination and diagnosis responses are correct, but the treatment response is incorrect. When only the relevant medical records are added, the examination response is incorrect, while diagnosis and treatment level responses are correct. Finally, when both the experience and records are added simultaneously, the doctor can provide correct responses for all three tasks.
[backgroundcolor=gray!8!white, linecolor=gray!80!white] Isabella Rodriguez: Can you tell me what symptoms or conditions you are experiencing, Klaus Mueller?
Klaus Mueller: Hi, doctor! I feel unwell with specific symptoms: Discomfort in the throat, Dry Cough, Dry Mouth. These symptoms have lasted for 2 weeks. What is the cause of my illness?
Isabella Rodriguez: You need to do the following examination: blood test
Klaus Mueller: I have done the physical examination you requested, and here is the medical examination report.
Isabella Rodriguez: Your medical examination result of blood test is: If a patient experiences bacterial infection, blood routine examination may reveal an increase in the total white blood cell count and/or the proportion of neutrophils.
Klaus Mueller: According to your examination results, your diagnosis is chronic bronchitis, and the condition is categorized as mild.
Isabella Rodriguez: Here are some treatment plans: Patients should actively quit smoking; general measures must include attention to nutrition, increased outdoor activities, and physical exercise. Prompt treatment should be given to related causes such as sinusitis and adenoiditis. Attention should be paid to seasonal changes and avoidance of potential allergens to reduce the frequency of attacks. Children with excessive phlegm can use guaifenesin, ambroxol, and mucinolytic agents (myrtol) as mucolytic expectorants, which also increase mucous membrane ciliary movement and help in the excretion of sputum.
Klaus Mueller: Thank you for your valuable advice and support. I truly appreciate it. Goodbye!
Isabella Rodriguez: Take care and follow the treatment plan closely. Goodbye and best wishes for your recovery!
@@ Appendix A Contributions of Authors
Appendix B Acknowledgments
This work is supported by the National Natural Science Foundation of China (No. 61925601, 62276152, 62372260) and AI Industry Research Innovation Center (AIRIC), Wuxi Research Institute for Applied Technologies, Tsinghua University.
Appendix C Involved Disease Knowledge
C.1. Diseases
@@ C.2. Data Distribution
C.3. An Example of electronic Health Records of a Patient